# DiBS-MTL: Transformation-Invariant Multitask Learning with Direction Oracles

### v1

@@ Abstract
Multitask learning (MTL) algorithms typically rely on schemes that combine different task losses or their gradients through weighted averaging. These methods aim to find Pareto stationary points by using heuristics that require access to task loss values, gradients, or both. In doing so, a central challenge arises because task losses can be arbitrarily, nonaffinely scaled relative to one another, causing certain tasks to dominate training and degrade overall performance. A recent advance in cooperative bargaining theory, the Direction-based Bargaining Solution (DiBS), yields Pareto stationary solutions immune to task domination because of its invariance to monotonic nonaffine task loss transformations. However, the convergence behavior of DiBS in nonconvex MTL settings is currently not understood. To this end, we prove that under standard assumptions, a subsequence of DiBS iterates converges to a Pareto stationary point when task losses are possibly nonconvex, and propose DiBS-MTL, a computationally efficient adaptation of DiBS to the MTL setting. Finally, we validate DiBS-MTL empirically on standard MTL benchmarks, showing that it achieves competitive performance with state-of-the-art methods while maintaining robustness to nonaffine monotonic transformations that significantly degrade the performance of existing approaches, including prior bargaining-inspired MTL methods. Code available at https://github.com/suryakmurthy/dibs-mtl
@@ 1 Introduction
The successes of deep learning have inspired investigation into “generalist” networks—models simultaneously trained for learning multiple tasks. As a result, numerous multitask learning (MTL) algorithms have been developed to tackle the inevitable conflict between task-specific loss gradients, aiming to ensure that during training, no task is under-optimized compared to others (Kendall et al., 2018; Sener & Koltun, 2018; Yu et al., 2020a; Liu et al., 2021a; Navon et al., 2022; Liu et al., 2023). However, most existing MTL methods are not robust against non-affine (monotonic) transformations to task losses, which is a crucial property desirable in the context of deep learning—where same preferences can be represented with losses of different nonaffine scalings, and it is unclear which relative scaling of the different losses ensures balanced learning without expensive and exhaustive ablations. We consider the problem of multitask learning (MTL) through the lens of centralized, cooperative bargaining methods that are invariant to non-affine monotonic task loss transformations.
The issue of different task losses being directly incomparable and scaled in different, non-affine fashions arises very naturally in many deep learning domains. For instance, reinforcement learning applications demand that a practitioner leverages prior, task-specific domain knowledge to design an effective reward function (Yu et al., 2025). However, in a downstream MTL setting, the loss corresponding to this reward function may dominate (or get dominated by) other task losses. At the same time, the relative performance of a “good” task-specific policy does not change when the corresponding task loss is monotonically transformed,i.e., the transformation does not change the actual underlying preferences over options. Thus, in an MTL problem, the available task losses can be seen as monotonic—possibly non-affine—transformations of some underlying set of ideal, unknown task losses that are meaningfully scaled with respect to each other.
Recent work in MTL has developed a connection with cooperative game theory (Navon et al., 2022). In this setting, each different loss function is a separate player in a bargaining game, and the idea is to find a balanced Pareto optimum among the players’ objectives. Classical solutions to these bargaining games (e.g., Nash, as explored by Navon et al. (2022)) are not robust to non-affine monotonic scalings, and only recently has a technique—Direction-based Bargaining Solution (DiBS)—been developed which remains invariant to these transformations (Gupta et al., 2025). However, the convergence of DiBS has only been analyzed in settings with strongly convex losses, and it is unclear to what extent the favorable properties of DiBS will apply in realistic MTL applications where task losses are almost always nonconvex. Inspired by this, we investigate the following:
@@
    What theoretical properties can be established for Direction-based Bargaining Solution (DiBS) in the general setting where player objectives (task losses) can be non-convex?
    Contribution 1. We show that under standard assumptions, for non-convex losses, a subsequence of the DiBS iterates provably converges to a Pareto stationary point asymptotically. Notably, our result does not require the linear independence of task loss gradients at non-Pareto stationary points, an assumption that is required by MTL methods using the Nash bargaining solution to deliver the same asymptotic guarantee (Navon et al., 2022).
@@
    Contribution 2. We extend DiBS to multitask learning, showing its natural compatibility with existing bargaining-for-learning frameworks. Moreover, we propose an approximation, DiBS-MTL, which is computationally more efficient than DiBS, and also preserves desirable invariance to non-affine monotonic task loss transformations. We empirically show that DiBS-MTL performs competitively with existing MTL methods on widely used multitask computer vision and reinforcement learning benchmarks.
    Contribution 3. We further investigate multitask reinforcement learning settings in which different task reward functions undergo non-affine monotonic transformations—causing potentially non-monotonic critic loss transformations. We empirically demonstrate that in this setting, DiBS-MTL maintains robust performance, while existing state-of-the-art MTL methods observe a significant drop in performance under the same transformations.
@@ 2.1 Related MTL literature
The most popular MTL approach in practice is linear scalarization (LS)—constructing a scalarized loss by taking the unweighted sum of task losses, or using known static coefficients to compute a weighted average. While previous work has advocated for LS methods (Kurin et al., 2022; Xin et al., 2022), in practice, LS can lead to situations where certain tasks remain under-optimized. Further, it has been shown that LS also does not necessarily recover the entire Pareto front generated by the task losses (Hu et al., 2023). Other methods tackle the MTL problem through more sophisticated multiobjective optimization approaches, aiming to find Pareto optimal (or stationary, in general nonconvex settings) points. Such methods seek to address the task imbalances arising during training via heuristics that (i) use task-specific loss values to compute a scalar weighted average loss (but with evolving weights, unlike LS) (Kendall et al., 2018; Liu et al., 2019; Lin et al., 2022; Liu et al., 2023), or (ii) use task-specific loss gradients to find update directions iteratively during training (Sener & Koltun, 2018; Yu et al., 2020a; Chen et al., 2020; Liu et al., 2021a; b; Navon et al., 2022).
@@
Existing loss-based heuristics include maximizing improvement of the worst-performing task (Liu et al., 2023), forcing improvements to be similar across tasks (Liu et al., 2019), weighting task losses randomly (Lin et al., 2022), and adapting weights according to task-based uncertainty measures (Kendall et al., 2018). Heuristics for gradient-based approaches include finding mutual task improvement directions (Yu et al., 2020a), probabilistically masking task gradients according to their sign (Chen et al., 2020), computing weights that minimize the norm of the convex combination of task gradients (Sener & Koltun, 2018), using gradient-based approaches to maximize improvement of the worst-performing task (Liu et al., 2021a), and finding the Nash bargaining solution for a bargaining subproblem (Navon et al., 2022). However, as we empirically demonstrate in Section 5, these existing methods are not robust to monotonic, nonaffine task loss transformations, potentially experiencing significant performance drops in such settings. We remark than one gradient-based method, IMTL-G (Liu et al., 2021b), in principle can produce solutions invariant to monotonic nonaffine task loss transformations. However, we show in Appendix C that even for simple convex problems, IMTL-G can converge to Pareto solutions that heavily favor one task over the other.
@@
Finally, we note that other approaches also exist for the MTL problem, such as (i) task clustering, where methods first cluster tasks to reduce conflicts, and then update model training parameters for each cluster (Standley et al., 2020; Fifty et al., 2021; Song et al., 2022; Shen et al., 2024); and (ii) parameter sharing methods, which design neural network architectures consisting of task-specific and task-shared modules/parameters (Kokkinos, 2017; Guo et al., 2020; Gao et al., 2020). These approaches differ from the proposed approach at a fundamental level, in that they try to design a suite of models for the space of tasks, or design model architectures suitable for MTL, whereas the proposed approach aims to optimize a a single model (with a pre-defined architecture), which balances the performance of all tasks.
@@ 2.2 Preliminaries on cooperative bargaining games
$\mathbf{x}_{k+1}=\mathbf{x}_{k}-h(\mathbf{x}_{k}):=\mathbf{x}_{k}-\alpha_{k}\cdot\sum_{i\in[N]}\left(\|\mathbf{x}_{k}-\mathbf{x}^{*,i}\|_{2}\cdot\frac{\nabla_{\mathbf{x}}\ell^{i}(\mathbf{x}_{k})}{\|\nabla_{\mathbf{x}}\ell^{i}(\mathbf{x}_{k})\|_{2}}\right).$ (1)
@@
We also note that prior has work has used the classical Nash bargaining solution (NBS) (Nash et al., 1950) in the context of MTL (Navon et al., 2022). However, the NBS is invariant to only affine agent (or task) loss transformations, and can change if the agent (or task) losses undergo monotonic nonaffine transformations.
As highlighted in Section 1, the need for MTL methods which are robust to monotonic non-affine transformations makes it natural to consider incorporating DiBS—which is invariant to such distortions—in an MTL approach. However, since practical MTL problems typically involve nonconvex losses, it is crucial to examine the behavior of DiBS in the nonconvex regime, a setting not yet analyzed in prior work. This serves as our motivation for the next section.
@@ 3 What can we say about DiBS in the nonconvex regime?
The sequence $\{\mathbf{x}_{k}\}_{k=1}^{\infty}$ generated by the DiBS iterations given in Equation 1 are bounded, i.e., there is a bounded set $\mathcal{D}\subseteq\mathcal{S}$ such that $\mathbf{x}_{k}\in\mathcal{D}~\forall~k\in\mathbb{N}$.
@@
Let $\{\mathbf{x}_{k}\}_{k=1}^{\infty}$ be the sequence generated by the DiBS algorithm given in Equation 1, for an initial point $\mathbf{x}_{1}\in\mathcal{S}$ and stepsizes that follow the Robbins-Monro conditions, i.e., $\sum_{k}\alpha_{k}=\infty$, $\sum_{k}\alpha_{k}^{2}<\infty$ (Robbins & Monro, 1951). Then, under Assumptions 1 and 2, the sequence $\{\mathbf{x}_{k}\}_{k=1}^{\infty}$ has a subsequence that asymptotically converges to a Pareto stationary point $\mathbf{x}^{\dagger}$, i.e., $h(\mathbf{x}^{\dagger})=0$.
@@
4 Adapting DiBS to multitask learning
We now extend DiBS to the multitask learning setting. We begin by introducing some notation, and then describe the underlying bargaining game for MTL, which has been proposed in previous work (Navon et al., 2022).
@@ 4 Adapting DiBS to multitask learning
We use $\bm{0}$ to denote a zero vector of appropriate dimensions, and $\mathcal{V}(\mathbf{x},r)$ denotes a ball of radius $r$, centered at $\mathbf{x}$. The vector $\theta\in\mathbb{R}^{n}$ represents shared task parameters; any additional task-specific parameters are suppressed for notational convenience and do not influence the calculation of $\theta$ iterates in this section.
@@
The goal in MTL is to train a model to perform multiple tasks. Every learning task has an associated task loss $\ell^{i}(\theta):\mathbb{R}^{n}\to\mathbb{R}$. During training, given parameters $\theta$, bargaining is iteratively conducted to find the next set of parameters $\theta+\Delta\theta$.
To this end, every task is represented as a distinct bargaining agent with the (minimization) objective being an approximation of the difference $\ell^{i}(\theta+\Delta\theta)-\ell^{i}(\theta)$. We consider a first-order approximation, in line with prior bargaining work for a fair comparison (Navon et al., 2022), with the $i^{\text{th}}$ agent’s objective being
$\min_{\|\Delta\theta\|_{2}\leq\epsilon}\underbrace{\left(\nabla_{\theta}\ell^{i}(\theta)\right)^{\top}\Delta\theta}_{:=\omega^{i}(\Delta\theta)},$
where we constrain the update vector $\Delta\theta$ to lie in $\mathcal{V}(\bm{0},\epsilon)$, to prevent overshooting caused by large update steps. Thus, for an MTL problem with $N$ tasks—the bargaining game becomes $\mathcal{B}_{\mathcal{V}(\bm{0},\epsilon)}([\omega^{1},\dots,\omega^{N}])$.
@@
$\frac{\nabla_{\Delta\theta}\omega^{i}}{\|\nabla_{\Delta\theta}\omega^{i}\|_{2}}=\frac{\nabla_{\theta}\ell^{i}(\theta)}{\|\nabla_{\theta}\ell^{i}(\theta)\|_{2}}$
only needs to be computed once at the starting of a bargaining game. Further, the quantity $\mathbf{x}^{*,i}$ in Equation 1 has a closed form solution $\mathbf{x}^{*,i}=-\left(\nicefrac{{\nabla_{\theta}\ell^{i}(\theta)}}{{\|\nabla_{\theta}\ell^{i}(\theta)\|_{2}}}\right)\cdot\epsilon$ for the linear objectives, i.e., the furthest point in the negative gradient direction, given that $\Delta\theta$ is constrained to lie in $\mathcal{V}(0,\epsilon)$. Thus, for an initial choice $\Delta\theta_{1}$ and stepsize sequence $\{\alpha_{k}\}$, the DiBS iterates for MTL become
$\Delta\theta_{k+1}=\Delta\theta_{k}-\alpha_{k}\cdot\sum_{i\in[N]}\left(\left\|\Delta\theta_{k}+\epsilon\cdot\frac{\nabla_{\theta}\ell^{i}(\theta)}{\|\nabla_{\theta}\ell^{i}(\theta)\|_{2}}\right\|_{2}\cdot\frac{\nabla_{\theta}\ell^{i}(\theta)}{\|\nabla_{\theta}\ell^{i}(\theta)\|_{2}}\right),k\in\mathbb{N}$ (Multi-step DiBS-MTL)
In practice, naively applying Multi-step DiBS-MTL can be slow. We propose the following approximation to DiBS, which is more computationally efficient and still preserves the desirable invariance to monotonic, nonaffine task loss transformations:
$\Delta\theta=-\epsilon\cdot\sum_{i\in[N]}\frac{\nabla_{\theta}\ell^{i}(\theta)}{\|\nabla_{\theta}\ell^{i}(\theta)\|_{2}}$ (DiBS-MTL)
DiBS-MTL can be viewed as an approximation of Multi-step DiBS-MTL, obtained by setting the initial $\Delta\theta_{1}=0$, running the system for a single step, i.e., only for $k=1$, and outputting $\Delta\theta_{2}$ as $\Delta\theta$. Such single step approximations are common for ensuring practical computational speedups in MTL methods where an iterative processes are involved, c.f. FAMO (Liu et al., 2023). As we will show in Section 5, DiBS-MTL offers competitive performance on MTL tasks, even with a single step approximation. We remark that unlike equation 1, the single step update rule DiBS-MTL that does not have explicit dependencies on $\mathbf{x}^{*,i}$ due to the first-order approximation combined with the ball constraint. In Appendix E we provide results on a demonstrative two dimensional nonconvex example showing that the single-step version performs similarly to the multi-step version despite being relatively simpler.

### v2

# Monotonic Transformation Invariant Multi-task Learning

@@ Abstract
Multi-task learning (MTL) algorithms typically rely on schemes that combine different task losses or their gradients through weighted averaging. These methods aim to find Pareto stationary points by using heuristics that require access to task loss values, gradients, or both. In doing so, a central challenge arises because task losses can be arbitrarily scaled relative to one another, causing certain tasks to dominate training and degrade overall performance. A recent advance in cooperative bargaining theory, the Direction-based Bargaining Solution (DiBS), yields Pareto stationary solutions immune to task domination because of its invariance to monotonic nonaffine task loss transformations. However, the convergence behavior of DiBS in nonconvex MTL settings is currently not understood. To this end, we prove that under standard assumptions, a subsequence of DiBS iterates converges to a Pareto stationary point when task losses are nonconvex, and propose DiBS-MTL, an adaptation of DiBS to the MTL setting which is more computationally efficient that prior bargaining-inspired MTL approaches. Finally, we empirically show that DiBS-MTL is competitive with leading MTL methods on standard benchmarks, and it drastically outperforms state-of-the-art baselines in multiple examples with poorly-scaled task losses, highlighting the importance of invariance to nonaffine monotonic transformations of the loss landscape. Code available at https://github.com/suryakmurthy/dibs-mtl
Machine Learning, ICML
@@ 1 Introduction
The successes of deep learning have inspired investigation into “generalist” networks—models simultaneously trained for learning multiple tasks. As a result, numerous multi-task learning (MTL) algorithms have been developed to tackle the inevitable conflict between task-specific loss gradients, aiming to ensure that during training, no task is under-optimized compared to others (Kendall et al., 2018; Sener and Koltun, 2018; Yu et al., 2020a; Liu et al., 2021a; Navon et al., 2022; Liu et al., 2023). However, most existing MTL methods are not robust against nonaffine (monotonic) transformations to task losses, which is a crucial property desirable in the context of deep learning—where same preferences can be represented with losses of different nonaffine scalings, and it is unclear which relative scaling of the different losses ensures balanced learning without expensive and exhaustive ablations. We consider the problem of multi-task learning (MTL) through the lens of centralized, cooperative bargaining methods that are invariant to nonaffine monotonic task loss transformations.
The issue of different task losses being directly incomparable and scaled in different, nonaffine fashions arises very naturally in many deep learning domains. For instance, reinforcement learning applications demand that a practitioner leverages prior, task-specific domain knowledge to design an effective reward function (Yu et al., 2025). However, in a downstream MTL setting, the loss corresponding to this reward function may dominate (or get dominated by) other task losses. At the same time, the relative performance of a “good” task-specific policy does not change when the corresponding task loss is monotonically transformed, i.e., the transformation does not change the actual underlying preferences over options. Thus, in an MTL problem, the available task losses can be seen as monotonic, possibly nonaffine, transformations of some underlying set of ideal, unknown task losses that are meaningfully scaled with respect to each other.
Recent work in MTL has developed a connection with cooperative game theory (Navon et al., 2022). In this setting, each different loss function is a separate agent in a bargaining game, and the idea is to find a balanced Pareto optimum among the agents’ objectives. Classical solutions to these bargaining games (e.g., Nash, as explored by Navon et al. (2022)) are not robust to nonaffine monotonic scalings, and only recently has a technique—Direction-based Bargaining Solution (DiBS)—been developed which remains invariant to these transformations (Gupta et al., 2025). However, the convergence of DiBS has only been analyzed in settings with strongly convex losses, and it is unclear to what extent the favorable properties of DiBS will apply in realistic MTL applications where task losses are almost always nonconvex. Inspired by this, we investigate the following:
@@
    What theoretical properties can be established for Direction-based Bargaining Solution (DiBS) in the general setting where agent objectives (task losses) can be nonconvex?
    Contribution 1. We show that under standard assumptions, for nonconvex losses, a subsequence of the DiBS iterates converges to a Pareto stationary point asymptotically. Notably, our result does not require the linear independence of task loss gradients at non-Pareto stationary points, an assumption that is required by MTL methods using the Nash bargaining solution to deliver the same asymptotic guarantee (Navon et al., 2022).
@@
    Contribution 2. We extend DiBS to multi-task learning, showing its natural compatibility with existing bargaining-for-learning frameworks. To this end, we propose DiBS-MTL, which adapts DiBS to MTL utilizing local loss approximations, is computationally more efficient than previous bargaining-inspired MTL approaches, and also preserves desirable invariance to nonaffine monotonic task loss transformations. We empirically show that DiBS-MTL performs competitively with state-of-the-art MTL methods on widely used vision and quantum chemistry MTL benchmarks.
    Contribution 3. We further investigate settings where task losses are nonaffinely and poorly scaled relative to each other in two domains—a challenging synthetic problem that exhibits substantial nonconvexity (and which highlights Contribution 1), and a larger-scale multi-task reinforcement learning benchmark. We show that DiBS-MTL drastically outperforms state-of-the-art baselines in these domains; highlighting the importance of MTL methods to be invariant to nonaffine monotonic transformations of the loss landscape.
@@ 2.1 Related MTL literature
The most popular MTL approach in practice is linear scalarization (LS)—constructing a scalarized loss by taking the unweighted sum of task losses, or using known static coefficients to compute a weighted average. While previous work has advocated for LS methods (Kurin et al., 2022; Xin et al., 2022), in practice, LS can lead to situations where certain tasks remain under-optimized. Further, it has been shown that LS also does not necessarily recover the entire Pareto front generated by the task losses (Hu et al., 2023). Other methods tackle the MTL problem through more sophisticated multiobjective optimization approaches, aiming to find Pareto optimal (or stationary, in general nonconvex settings) points. Such methods seek to address the task imbalances arising during training via heuristics that (i) use task-specific loss values to compute a scalar weighted average loss (but with evolving weights, unlike LS) (Kendall et al., 2018; Liu et al., 2019; Lin et al., 2022; Liu et al., 2023), or (ii) use task-specific loss gradients to find update directions iteratively during training (Sener and Koltun, 2018; Yu et al., 2020a; Chen et al., 2020; Liu et al., 2021a, b; Navon et al., 2022).
@@
Existing loss-based heuristics include maximizing improvement of the worst-performing task (Liu et al., 2023), forcing improvements to be similar across tasks (Liu et al., 2019), weighting task losses randomly (Lin et al., 2022), weighting task losses according to relative improvements (Chen et al., 2018), and adapting weights according to task-based uncertainty measures (Kendall et al., 2018). Heuristics for gradient-based approaches include finding mutual task improvement directions (Yu et al., 2020a), probabilistically masking task gradients according to their sign (Chen et al., 2020), computing weights that minimize the norm of the convex combination of task gradients (Sener and Koltun, 2018), using gradient-based approaches to maximize improvement of the worst-performing task (Liu et al., 2021a), and finding the Nash bargaining solution for a bargaining subproblem (Navon et al., 2022). However, as we empirically demonstrate in Section 5, these existing methods are not robust to monotonic, nonaffine task loss transformations, potentially experiencing significant performance drops in such settings. We remark than one gradient-based method, IMTL-G (Liu et al., 2021b), in principle can produce solutions invariant to monotonic nonaffine task loss transformations. However, we show in Appendix C that even for simple convex problems, IMTL-G can converge to Pareto solutions that heavily favor one task over the other.
@@
Finally, we note that other approaches also exist for the MTL problem, such as (i) task clustering, where methods first cluster tasks to reduce conflicts, and then update model training parameters for each cluster (Standley et al., 2020; Fifty et al., 2021; Song et al., 2022; Shen et al., 2024); and (ii) parameter sharing methods, which design neural network architectures consisting of task-specific and task-shared modules/parameters (Kokkinos, 2017; Guo et al., 2020; Gao et al., 2020). These approaches differ from the proposed approach at a fundamental level, in that they try to design a suite of models for the space of tasks, or design MTL model architectures, whereas the proposed approach aims to optimize a single model (with a pre-defined architecture), balancing the performance of all tasks.
@@ 2.2 Preliminaries on cooperative bargaining games
$\mathbf{x}_{k+1}=\mathbf{x}_{k}-h(\mathbf{x}_{k})$
$:=$ $\mathbf{x}_{k}-\alpha_{k}\cdot\sum_{i\in[N]}\left(\|\mathbf{x}_{k}-\mathbf{x}^{*,i}\|_{2}\cdot\frac{\nabla_{\mathbf{x}}\ell^{i}(\mathbf{x}_{k})}{\|\nabla_{\mathbf{x}}\ell^{i}(\mathbf{x}_{k})\|_{2}}\right).$ (1)
@@
We also note that prior has work has used the classical Nash bargaining solution (NBS) (Nash and others, 1950) in the context of MTL (Navon et al., 2022). However, the NBS is invariant to only affine agent (or task) loss transformations, and can change if the agent (or task) losses undergo monotonic nonaffine transformations.
As highlighted in Section 1, the need for MTL methods which are robust to monotonic nonaffine transformations makes it natural to consider incorporating DiBS—which is invariant to such distortions—in an MTL approach. However, since practical MTL problems typically involve nonconvex losses, it is crucial to examine the behavior of DiBS in the nonconvex regime, a setting not yet analyzed in prior work. This serves as our motivation for the next section.
@@ 3 What can we say about DiBS in the nonconvex regime?
The sequence $\{\mathbf{x}_{k}\}_{k=1}^{\infty}$ generated by the DiBS iterations given in Section 2.2 are bounded, i.e., there is a bounded set $\mathcal{D}\subseteq\mathcal{S}$ such that $\mathbf{x}_{k}\in\mathcal{D}~\forall~k\in\mathbb{N}$.
@@
Let $\{\mathbf{x}_{k}\}_{k=1}^{\infty}$ be the sequence generated by the DiBS algorithm given in Section 2.2, for an initial point $\mathbf{x}_{1}\in\mathcal{S}$ and stepsizes that follow the Robbins-Monro conditions, i.e., $\sum_{k}\alpha_{k}=\infty$, $\sum_{k}\alpha_{k}^{2}<\infty$ (Robbins and Monro, 1951). Then, under Assumptions 1 and 2, the sequence $\{\mathbf{x}_{k}\}_{k=1}^{\infty}$ has a subsequence that asymptotically converges to a Pareto stationary point $\mathbf{x}^{\dagger}$, i.e., $h(\mathbf{x}^{\dagger})=0$.
@@
Algorithm 1 $T$-step DiBS-MTL
0: Initial parameters $\theta_{0}$, losses $\{\ell^{i}\}_{i=1}^{N}$, learning rate $\eta$, approximation radius $\epsilon$, step size $\alpha<\nicefrac{{\epsilon}}{{N}}$, number of epochs $J$
for $j=1$ to $J$ do
  Compute normalized gradients $\bar{g}^{i}_{j-1}=\frac{\nabla_{\theta}\ell^{i}(\theta_{j-1})}{\|\nabla_{\theta}\ell^{i}(\theta_{j-1})\|_{2}}$
  For each $i$, compute preferred individual parameters inside $\epsilon$-radius: $\theta^{*,i}_{j}=\theta_{j-1}-\epsilon\bar{g}^{i}_{j-1},$
  Set $\theta_{(j-1,0)}\leftarrow\theta_{j-1}$
  // DiBS updates inside approximation radius $\epsilon$
  for $t=1$ to $T$ do
   Compute DiBS update direction: $d_{(j,t)}=\sum_{i\in[N]}\Big(\|\theta_{(j-1,t-1)}-\theta^{*,i}_{j}\|_{2}\cdot\bar{g}^{i}_{j-1}\Big)$
   Update: $\theta_{(j-1,t)}\leftarrow\theta_{(j-1,t-1)}-\alpha\,d_{(j,t)}$
  end for
  $\Delta\theta_{j-1}=\theta_{(j-1,T)}-\theta_{(j-1,0)}$
  Update: $\theta_{j}\leftarrow\theta_{j-1}+\eta\,\Delta\theta_{j-1}$
end for
return $\theta_{j}$
4 Adapting DiBS to multi-task learning
We now extend DiBS to the multi-task learning setting. We begin by introducing some notation, and then describe the underlying bargaining game for MTL, which has been proposed in previous work (Navon et al., 2022).
@@ 4 Adapting DiBS to multi-task learning
We use $\bm{0}$ to denote a zero vector of appropriate dimensions, and $\mathcal{V}(\mathbf{x},r)$ denotes a ball of radius $r$, centered at $\mathbf{x}$. The vector $\theta\in\mathbb{R}^{n}$ represents shared task parameters; any additional task-specific parameters are suppressed for notational convenience and do not influence the calculation of $\theta$ iterates in this section. $j$ denotes an epoch iteration, whereas $t$ will denote bargaining updates within an epoch.
@@
The goal in MTL is to train a model to perform multiple tasks. Every learning task has an associated task loss $\ell^{i}(\theta):\mathbb{R}^{n}\to\mathbb{R}$. During training, given parameters $\theta_{j}$ at some epoch $j+1$, bargaining is iteratively conducted during the epoch to find the next set of parameters $\theta_{j+1}=\theta_{j}+\eta\Delta\theta_{j}$ for some learning rate $\eta$.
To this end, every task is represented as a distinct bargaining agent with the (minimization) objective being an approximation of the difference $\ell^{i}(\theta_{j}+\eta\Delta\theta_{j})-\ell^{i}(\theta_{j})$. We consider a first-order approximation, in line with prior bargaining work for a fair comparison (Navon et al., 2022), with the $i^{\text{th}}$ agent’s objective being
$\min_{\|\Delta\theta_{j}\|_{2}\leq\epsilon}\underbrace{\left(\nabla_{\theta}\ell^{i}(\theta_{j})\right)^{\top}\Delta\theta_{j}}_{:=\omega^{i}(\Delta\theta_{j})},$
where we constrain the update vector $\Delta\theta_{j}$ to lie in $\mathcal{V}(\bm{0},\epsilon)$, to prevent overshooting caused by large update steps. Thus, for an MTL problem with $N$ tasks—the bargaining game becomes $\mathcal{B}_{\mathcal{V}(\bm{0},\epsilon)}([\omega^{1},\dots,\omega^{N}])$.
@@
$\frac{\nabla_{\Delta\theta_{j}}\omega^{i}}{\|\nabla_{\Delta\theta_{j}}\omega^{i}\|_{2}}=\frac{\nabla_{\theta}\ell^{i}(\theta_{j})}{\|\nabla_{\theta}\ell^{i}(\theta_{j})\|_{2}}:=\bar{g}^{i}_{j}$
only needs to be computed once at the starting of a bargaining game. Further, the distance of the individual optima from the initial point in Section 2.2 has a closed form solution such that $\Delta\theta^{*,i}_{j}:=\theta^{*,i}_{j}-\theta_{j-1}=-\left(\nicefrac{{\nabla_{\theta}\ell^{i}(\theta_{j})}}{{\|\nabla_{\theta}\ell^{i}(\theta_{j})\|_{2}}}\right)\cdot\epsilon$ for the linear objectives, i.e., an individual optimum is the furthest point in the negative gradient direction, given that $\Delta\theta_{j}$ is constrained to lie in $\mathcal{V}(\bm{0},\epsilon)$. Thus, for an initial choice $\Delta\theta_{(j,0)}=\bm{0}$ and stepsize sequence $\{\alpha_{t}\}$, $T$-iterate DiBS  for MTL satisfies
$\Delta\theta_{(j,t+1)}=\Delta\theta_{(j,t)}-\alpha_{t}\cdot\sum_{i\in[N]}\Bigg(\left\|\Delta\theta_{(j,t)}+\epsilon\cdot\bar{g}^{i}_{j}\right\|_{2}\cdot\bar{g}^{i}_{j}\Bigg)$ (T-step DiBS-MTL)
Using this the parameters between epochs are updated as in Algorithm 1. In practice, single-step versions of such iterative schemes are common (Liu et al., 2023). Such a version described below still preserves the desirable invariance to monotonic, nonaffine task loss transformations:
$\Delta\theta_{(j,1)}=-\epsilon\cdot\sum_{i\in[N]}\bar{g}^{i}_{j}$ (1-step DiBS-MTL)
We remark that unlike multi-step variants, 1-step DiBS-MTL does not have explicit dependencies on the individual optima $\theta^{*,i}_{j}$ due to the first-order approximation combined with the ball constraint. Further, 1-step DiBS-MTL does not introduce any additional hyperparameters as the $\epsilon$ term is absorbed into the learning rate $\eta$. In Sections 5.1, 5.3 and F we provide results showing that the single-step version performs similarly to the multi-step version despite being relatively simpler on a variety of MTL problems.